Consider the reaction for the preparation of the aluminum phosphate precipitate:
Na3PO4(aq)+ALCL3(aq) —> ALPO4(s)+3BaCL(aq)
If 20g of sodium phosphate react with an equal mass of aluminium chloride calculate:
A) number of moles of the precipitate formed
B)mass of remaining excess reactant
C)actual yield of the reaction if percentage yield is 80%
1
Expert's answer
2019-04-23T01:59:41-0400

"n(Na3PO4)=\dfrac{20g}{163.94g/mol}=0.122 moles"

"n(AlCl3)=\dfrac{20g}{133.34g/mol}=0.15mol"


"n(remaining.excess.reactant(AlCl3))=0.15moles - 0.122moles=0.028moles"

"m(remaining.excess.reactant(AlCl3))=133.34g\/mol*0.028moles=3.734g"


"m(AlPO4)=\dfrac{m(Na3PO4)*M(AlPO4)}{M(Na3PO4)}=\dfrac{20g*121.95g/mol}{163.94g/mol}=14.87g"

Actual yield"=14.87*80\/100=11.896g"

Answer:

A)0.122moles

B)3.734g

C)11.896g